I am having trouble finding the length and width if a certain rectangle has an area of 288 square inches and a perimeter of 9 feet.
I changed 9 feet into inches which is 108 inches.
I tried to do guess and check, but it is not really working for me...

Found 2 solutions by ewatrrr, solver91311:
Answer by ewatrrr(24785)

 
Hi,
P(rectangle) = 2L + 2W
2L + 2W = 108in
L + W = 54
L = (54-W)
(54-W)W = 288 in^2
W^2 - 54W + 288 = 0
factoring
(W -48)(W-6) = 0
(W -48)= 0
(W-6) = 0
rectangle is 6in by 48in or 48in by 6in

Answer by solver91311(24713)



The perimeter of a rectangle is given by:



and the area is given by:



Plug in numbers:



which means



Then



But since we have an expression for the length in terms of width that we derived from the Area equation, substitute:



Multiply by



Put in standard form:



Factor and solve for . Then calculate


John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism